OpenStudy (anonymous):

find the length of the side of a square with an area of 121 484 30.25 11 6.5

OpenStudy (anonymous):

Do you know the formula for the area of a square?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

l*w, what number multiplied by itself equals 121?

OpenStudy (anonymous):

60.5

OpenStudy (nurali):

\[Area=a^2\] \[a=\sqrt{A}\] \[a =\sqrt{121}\]

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Yes, 11*11 equals 121

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No, you were right when you said the answer was 11

OpenStudy (anonymous):

i know my computer is acting crazy

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you feel like helping me with another one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

r u a teacher?

OpenStudy (anonymous):

which is the closest to the surface area in a square inches that jim is going to refinish 19.5 inches answers are 1200 1600 2000 2400

OpenStudy (anonymous):

no not a teacher

OpenStudy (anonymous):

So is the surface area of a square, and the length of one side is 19.5 inches?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To find the S.A. of a cube, you would find the area of one side, and then multiply by 6. 19.5*19.5=? ?*6=?

OpenStudy (anonymous):

2281.50 is what i got...ughhh

OpenStudy (anonymous):

Yeah, that's the answer for the surface area of a cube that has the side length of 19.5 inches. Why isn't it an answer choice?

OpenStudy (anonymous):

19.5x2=380.25

OpenStudy (anonymous):

380.25x4=1521

OpenStudy (anonymous):

so would 1600 be the answer

OpenStudy (anonymous):

It doesn't make sense. Are you sure that we need the surface area of a cube?

OpenStudy (anonymous):

no, because a cube has 6 sides.

OpenStudy (anonymous):

it says old storage chest shaped like a cube

OpenStudy (anonymous):

Let's try rounding 19.5 to 20. 20*20=? ?*6=?

OpenStudy (anonymous):

2400 ?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

That's the answer I would go with.

OpenStudy (anonymous):

you up for one more???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Find the volume of a triangular prism with the flowing demenisons. Round to the nearest whole number Triangle Base 7ft Triangle height 4ft prism height 8.5 answers 20 ft3 65 ft3 119 ft3 238 ft3

OpenStudy (anonymous):

r u there orange master

OpenStudy (nurali):

V = Area of base x height A = 1/2 (b)(h1) A = ½(7)(4) A = ½(35) A = 14 V = (14)(8.5) V = 119 ft^3